if 32 is added to 80%of a no the result is the no itself find the no

Answer by ewatrrr(24785)

32 + .8n = n
32 = 1.0n - .8n
32 = .2n
32 = (2/10)n
320/2 = n
160 = n
And,..checking
32 + .8(160) = 32 + 128 = 160